The American Heart Association is about to conduct an anti-smoking campaign and wants to know the fraction of Americans over 20 who smoke. Step 1 of 2 : Suppose a sample of 966 Americans over 20 is drawn. Of these people, 783 don't smoke. Using the data, estimate the proportion of Americans over 20 who smoke. Enter your answer as a fraction or a decimal number rounded to three decimal places.
Answer #1

Solution: Americans over 20 is drawn sample = 966,

don't smoke = 783

smoke = 966 - 783 = 183

The proportion of Americans over 20 smoke is 0.189

=> p = X/n = 183/966 = 0.1894